# references.bib

### v2

@@ Appendix A Overview of Finnish Anti-Discrimination Law
Person A reported the case to the Non-Discrimination Ombudsman (Yhdenvertaisuusvaltuutettu), who brought the case before the National Non-Discrimination and Equality Tribunal (Yhdenvertaisuus-ja tasa-arvolautakunta).
Section 8(1) of the Non-Discrimination Act (FinlandNDA) defines the protected characteristics as: [^2] Official English translation, although only legally binding in Swedish (not included) and Finnish: “Syrjinnän kielto Ketään ei saa syrjiä iän, alkuperän, kansalaisuuden, kielen, uskonnon, vakaumuksen, mielipiteen, poliittisen toiminnan, ammattiyhdistystoiminnan, perhesuhteiden, terveydentilan, vammaisuuden, seksuaalisen suuntautumisen tai muun henkilöön liittyvän syyn perusteella. Syrjintä on kielletty riippumatta siitä, perustuuko se henkilöä itseään vai jotakuta toista koskevaan tosiseikkaan tai oletukseen”
@@
Section 3(1) of the Non-Discrimination Act provides that: “Provisions on prohibition of discrimination based on gender and the promotion of gender equality are laid down in the Act on Equality between Women and Men (609/1986).” The Non-Discrimination Act can be applied in cases of multiple discrimination, even if gender is one of the grounds of discrimination (FinlandGender; FinlandNDA, s 3(1)).
It is worth noting that this definition is broader than in the UK Equality Act. Some protected characteristics are outlined more explicitly; for example, a person discriminated against on the basis of language may be able to bring a claim based on racial discrimination (dziedziak2012). Unlike many Nordic countries, the Equality Act does not explicitly protect political activity, trade union activity, and does not include “or other personal characteristics” (hellum2023).
Direct discrimination is defined in Section 10:333“Syrjintä on välitöntä, jos jotakuta kohdellaan henkilöön liittyvän syyn perusteella epäsuotuisammin kuin jotakuta muuta on kohdeltu, kohdellaan tai kohdeltaisiin vertailukelpoisessa tilanteessa.”
@@
Indirect discrimination is defined in Section 13:444“Syrjintä on välillistä, jos näennäisesti yhdenvertainen sääntö, peruste tai käytäntö saattaa jonkun muita epäedullisempaan asemaan henkilöön liittyvän syyn perusteella, paitsi jos säännöllä, perusteella tai käytännöllä on hyväksyttävä tavoite ja tavoitteen saavuttamiseksi käytetyt keinot ovat asianmukaisia ja tarpeellisia.”
@@
Section 11(1) defines justifications for different treatment as:555“Erilainen kohtelu ei ole syrjintää, jos kohtelu perustuu lakiin ja sillä muutoin on hyväksyttävä tavoite ja keinot tavoitteen saavuttamiseksi ovat oikeasuhtaisia.”

### v3

@@ Appendix A Overview of Finnish Anti-Discrimination Law
Person A reported their case to the Non-Discrimination Ombudsman (Yhdenvertaisuusvaltuutettu), who brought the case before the National Non-Discrimination and Equality Tribunal (Yhdenvertaisuus-ja tasa-arvolautakunta). The Finnish Non-Discrimination Act is the relevant law (Finland Ministry of Justice, 2014b). Important extracts are quoted here using the official English translation, although only the Finnish and Swedish (not included) language is legally binding.
Section 8(1) of the Non-Discrimination Act defines the protected characteristics as:
@@
Syrjinnän kielto Ketään ei saa syrjiä iän, alkuperän, kansalaisuuden, kielen, uskonnon, vakaumuksen, mielipiteen, poliittisen toiminnan, ammattiyhdistystoiminnan, perhesuhteiden, terveydentilan, vammaisuuden, seksuaalisen suuntautumisen tai muun henkilöön liittyvän syyn perusteella. Syrjintä on kielletty riippumatta siitä, perustuuko se henkilöä itseään vai jotakuta toista koskevaan tosiseikkaan tai oletukseen.
Section 3(1) of the Non-Discrimination Act provides that: “Provisions on prohibition of discrimination based on gender and the promotion of gender equality are laid down in the Act on Equality between Women and Men (609/1986).” The Non-Discrimination Act can be applied in cases of multiple discrimination, even if gender is one of the grounds of discrimination (Finland Ministry of Justice, 2014a, b, s 3(1)).
It is worth noting that this definition is broader than in the UK Equality Act. Some protected characteristics are outlined more explicitly; for example, a person discriminated against on the basis of language may be able to bring a claim based on racial discrimination (United Kingdom Employment Appeals Tribunal, 2012). Unlike many Nordic countries, the UK Equality Act does not explicitly protect political activity, trade union activity, and does not include “or other personal characteristics” (Hellum et al., 2023).
Direct discrimination is defined in Section 10:
@@
Syrjintä on välitöntä, jos jotakuta kohdellaan henkilöön liittyvän syyn perusteella epäsuotuisammin kuin jotakuta muuta on kohdeltu, kohdellaan tai kohdeltaisiin vertailukelpoisessa tilanteessa.
Indirect discrimination is defined in Section 13:
@@
Syrjintä on välillistä, jos näennäisesti yhdenvertainen sääntö, peruste tai käytäntö saattaa jonkun muita epäedullisempaan asemaan henkilöön liittyvän syyn perusteella, paitsi jos säännöllä, perusteella tai käytännöllä on hyväksyttävä tavoite ja tavoitteen saavuttamiseksi käytetyt keinot ovat asianmukaisia ja tarpeellisia.
Section 11(1) defines justifications for different treatment as:
@@
Erilainen kohtelu ei ole syrjintää, jos kohtelu perustuu lakiin ja sillä muutoin on hyväksyttävä tavoite ja keinot tavoitteen saavuttamiseksi ovat oikeasuhtaisia.